OpenStudy (anonymous):

A ball is thrown vertically into the air from the ground with an initial velocity of 96 ft/sec. Its height, in feet, as a function of time, in seconds, is given by: h(t)=96t-16t 2 . Estimate the ball’s instantaneous velocity at time t = 3, including units. Explain what this value means in terms of the ball’s path. Instantaneous velocity at time t = 3 is: Intervals used and corresponding velocities: ???

OpenStudy (anonymous):

let's use v for velocity, h is for the distance in the vertical plane. we know \[v=\frac{ h }{ t }\] correct?

OpenStudy (anonymous):

OK...yeah

OpenStudy (anonymous):

Now, how do we find the h at t=3?

OpenStudy (anonymous):

I've been evaluating intervals into function at like 3.0 and 2.9?

OpenStudy (anonymous):

Is there a different way?

OpenStudy (anonymous):

what did you get for value of your distance?

OpenStudy (anonymous):

Like 1.6 ft/sec @ [2.9,3] & .016 ft/sec @ [2.999,3]

OpenStudy (anonymous):

From left to right it appears to be nearing zero?

OpenStudy (anonymous):

h(3) = 96(3)- 16 (3*3) = 144 ft/sec.

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Oh....is that it?

OpenStudy (anonymous):

I was taking the two intervals and dividing by the differnece. Getting zero

OpenStudy (anonymous):

Like....F9x1_-f(x0)/x1-x0

OpenStudy (anonymous):

Like....Fx1_-f(x0)/x1-x0

OpenStudy (anonymous):

use \[v = \frac{ h }{ t }\] Where h = 144 ft and t = 3 sec.

OpenStudy (anonymous):

is the question from physics or introductory calculus?

OpenStudy (anonymous):

Intro calc